# Optimizing Task Scheduling in Fog Computing with Deadline Awareness

### v4

@@ I Introduction
To maintain efficient resource allocation and establish QoS rules in IoT-fog networks, monitoring and analyzing network traffic is crucial. Traffic management becomes increasingly important as the number and complexity of Internet applications grow.
@@
This work proposes the Reinforcement Improved Golden Eagle Optimization (RIGEO) algorithm, an adaptive task scheduling framework. RIGEO dynamically classifies Fog Nodes (FNs) into low-traffic and high-traffic categories based on real-time network conditions. Low-traffic nodes are characterized by average network traffic that falls below a predefined threshold, whereas high-traffic nodes exceed this average. It employs Improved Golden Eagle Optimization (IGEO) for short-deadline tasks on low-traffic nodes and Reinforcement Learning for long-deadline tasks on high-traffic nodes, leveraging the strengths of each method for different operational scenarios.
RIGEO operates within a three-layered architecture: cloud, fog, and IoT devices. Figure 1 illustrates this hierarchy, where the fog layer comprises a controller and a network for resource management and task scheduling, while the cloud offers significant computing and storage capabilities. Since task scheduling is NP-hard [10], machine learning and heuristic methods are essential to achieve optimal performance [11].
The feasibility advantage of RIGEO over approaches that apply a single meta-heuristic to all tasks stems from matching algorithmic cost to task urgency. Population-based optimizers such as IGEO evaluate $N_{pop}\times N_{iter}$ candidate solutions per scheduling round, an overhead that is justified for short-deadline tasks, where even marginal improvements directly prevent costly deadline violations. For long-deadline tasks, however, this computational investment is excessive: a lightweight RL policy selects fog node assignments in $\mathcal{O}(|F_{\text{high}}|)$ per task by following a continuously updated transition-probability table, achieving competitive scheduling quality at negligible overhead. This selective application makes RIGEO practically deployable on resource-constrained fog controllers and scalable to the high task volumes characteristic of dense IoT deployments, where uniformly applying a meta-heuristic would saturate the scheduler before tasks can be dispatched.
@@ I-A Energy Efficiency
Vispute et al. [12] propose a PSO-based task scheduling scheme that outperforms existing approaches in makespan, energy consumption, and execution time, though performance may vary across scenarios with different resource limitations.
@@ I-B Deadline Constraints
Azami et al. [1] incorporate two semi-greedy algorithms to minimize deadline violations for IoT tasks while ensuring QoS compliance, achieving a significant reduction in deadline violation time, though performance may vary under different real-world conditions.
@@ I-C Latency Reduction
Klatoun et al. [7] propose a task scheduling framework for IoMT that processes data closer to end-users, reducing latency and enhancing response times. However, the study focuses on specific scenarios, limiting broader applicability.
Hosseini et al. [6] present an algorithm using fuzzy logic and AHP to prioritize tasks based on completion time, energy, RAM, and deadlines. While effective at balancing cost and latency in mobile fog environments, its complexity could challenge resource-constrained deployments.
I-D Reliability and Fault Tolerance
Ghanavati et al. [5] propose a learning automata-based task scheduling approach that ensures reliable execution while optimizing response time and energy consumption, though its focus on specific scenarios may limit broader applicability.
@@ I-D Reliability and Fault Tolerance
Energy Deadline Latency Reliability
@@
The remainder of this paper details the methodology in Section II, presents optimization results in Section III, and concludes in Section IV.
@@ II-A Problem Statement
The problem is to find an optimal mapping $M:T\rightarrow F$ to achieve these objectives.
Each task $t_{i}$ is described by a tuple $(c_{i},m_{i},d_{i},a_{i},q_{i})$: CPU demand $c_{i}\in\{1,\ldots,5\}$ cycles, normalized memory demand $m_{i}\in[0,1]$, deadline $d_{i}\in[0,1]$, discrete arrival time-slot $a_{i}\in\{1,\ldots,T\}$ where $T=10$, and minimum required QoS level $q_{i}\in[0,1]$. Tasks arrive according to a uniform discrete model: each task independently draws its arrival slot $a_{i}\sim\mathcal{U}\{1,\ldots,T\}$, so batch sizes vary naturally across slots.
Each FN $f_{j}$ is described by a tuple $(C_{j},M_{j},\kappa_{j},\theta_{j},\varepsilon_{j},h_{j})$: CPU capacity $C_{j}\in[81,100]$ units, memory capacity $M_{j}\in[5,16]$ GB, cost rate $\kappa_{j}\in[0,1]$ per cycle, QoS throughput $\theta_{j}\in[0.5,1.5]$, energy consumption rate $\varepsilon_{j}\in[3,6]$ J/cycle, and energy harvest rate $h_{j}=0.003$. All FN parameters are drawn independently and uniformly from these ranges, yielding a heterogeneous network. The fog network is modeled as a fully connected graph: every FN pair $(f_{i},f_{j})$ shares a symmetric communication-preference weight $P(i,j)=P(j,i)\sim\mathcal{U}[0,1]$, which the RL agent refines at runtime.
@@ II-A1 Response time
The response time for each task $t_{i}$ is determined by the sum of the propagation delay, transmission delay, queue waiting time, and execution time as shown in Eq. (1[9].
$R_{i}=P_{i}+T_{i}+Q_{i}+E_{i}$ (1)
Propagation delay, $P_{i}$, is the duration for a signal to move between locations within a communication channel. Transmission delay, $T_{i}$, is the time needed to transmit all task bits. Queue waiting time, $Q_{i}$, is the queue waiting time after task assignment, pending completion of higher-priority tasks. Execution time, $E_{i}$ is the CPU processing time to process the task.
@@ II-A2 Deadline violation
When assessing scheduling efficiency, it is essential to consider how many IoT tasks meet their deadlines compared to the total time spent on deadline violations. A deadline violation occurs when a task’s completion time exceeds its specified deadline [4].
@@ II-B Simulation Environment
The simulation models a fog layer of $m=20$ heterogeneous FNs representing edge servers with diverse computational capabilities, as summarized in Table II. All FNs are fully interconnected (all-to-all mesh), and communication preferences between any pair of nodes are captured by the symmetric probability matrix $P$ described above. No FN represents a fixed real-world device; instead, the heterogeneous parameter distributions replicate the diversity found in practical deployments of commodity edge hardware.
Five task-load scenarios are evaluated (200, 300, 400, 500, and 600 tasks) over 50 independent runs with different random seeds for task generation and node initialization; reported values are averages across these runs. Tasks arrive over $T=10$ discrete time slots; within each slot, all arriving tasks are dispatched together as a batch. Meta-heuristic algorithms (IGEO and all baselines) are allocated $N_{pop}=50$ candidate solutions and $N_{iter}=2$ optimization iterations per batch. Traffic patterns for all 20 FNs are simulated over 200 historical time steps, with each node’s load drawn uniformly from $[1,100]$ units; FNs whose mean traffic exceeds the network-wide median are labeled high-traffic ($F_{\text{high}}$), the remainder low-traffic ($F_{\text{low}}$). The deadline threshold separating short- and long-deadline tasks is fixed at the 25th percentile of each incoming batch’s deadline distribution, and the traffic threshold is recomputed as the median at the start of each run.
@@
Network topology Full mesh (all-to-all)
@@
FN CPU capacity ($C_{j}$) $\mathcal{U}[81,100]$ units
FN memory ($M_{j}$) $\mathcal{U}[5,16]$ GB
FN energy rate ($\varepsilon_{j}$) $\mathcal{U}[3,6]$ J/cycle
@@
FN cost rate ($\kappa_{j}$) $\mathcal{U}[0,1]$
@@
Deadline threshold 25th percentile of batch
Traffic threshold Median of node mean traffic
@@
II-C Proposed Method (RIGEO)
This work proposes RIGEO, a hybrid optimization approach that dynamically adapts to network conditions by classifying nodes based on traffic levels and applying different scheduling strategies based on task deadline requirements.
This work divides the network into two classes: low- and high-traffic. FNs are classified based on the median of observed network traffic. Nodes with traffic below the median are classified as low-traffic; otherwise, they are classified as high-traffic. Similarly, tasks are classified using the 25th percentile of the deadline distribution as the threshold, distinguishing time-critical tasks from those with relaxed time constraints. These thresholds are periodically recalculated to adapt dynamically to changing conditions.
Although periodic re-evaluation allows adaptation to evolving traffic patterns, handling sudden traffic spikes remains a challenge for future investigation. When network traffic is high, faster algorithms must be used to ensure that tasks are executed within their deadlines.
@@ II-C1 Reinforcement Learning for Long-Deadline Tasks
Reward and penalty. After task $t_{i}$ is assigned to FN $a$, the response time $R_{i}=c_{i}/(C_{a}+1)$ is compared with deadline $d_{i}$. A met deadline rewards the transition (scales up by 10%); a violation penalises it (scales down by 10%):
@@
This formulation limits per-task overhead to $\mathcal{O}(|F_{\text{high}}|)$, compared with $\mathcal{O}(N_{pop}\times N_{iter})$ for population-based methods, making it well-suited to long-deadline tasks whose relaxed timing constraint tolerates an incrementally improving policy rather than an exhaustive global search.
@@ II-C2 Golden Eagle Optimization (GEO)
GEO is a swarm-based meta-heuristic inspired by the hunting behavior of golden eagles. GEO’s search mechanism balances exploration (searching new areas) and exploitation (refining solutions). It adjusts speed during different hunting stages to find global optima while avoiding local optima.
In GEO, each eagle selects a target, calculates its cruising vector, and circles the best location found. The prey represents the optimal solution, with each eagle tracking its own best discovery. The attack vector for eagle $i$ is calculated through Eq. (8).
@@ II-C3 Improved Golden Eagle Optimization (IGEO)
Figure 2: Mutation operator
@@
Figure 3: Crossover operation. (a) One-point crossover. (b) Two-point crossover
@@
0: Task $T$ with deadline $D$, Set of Fog Nodes $F=\{f_{1},f_{2},\ldots,f_{n}\}$, traffic_thresholdtraffic\_threshold (median), deadline_thresholddeadline\_threshold (25th percentile)
@@
[…]
3:for each $f\in F$ do
4:  if f.traffic_level<traffic_thresholdf.traffic\_level<traffic\_threshold then
5:   Add $f$ to Low_Traffic_FNsLow\_Traffic\_FNs
6:  else
7:   Add $f$ to High_Traffic_FNsHigh\_Traffic\_FNs
8:  end if
9:end for
10:if T.deadline<deadline_thresholdT.deadline<deadline\_threshold then
11:  $f\leftarrow SELECT(Low\_Traffic\_FNs)$
12:  $result\leftarrow IGEO\_PROCESS(T,f)$
13:else
14:  $f\leftarrow SELECT(High\_Traffic\_FNs)$
15:  $result\leftarrow RL\_PROCESS(T,f)$
16:end if
17:return resultresult
(a) Energy Consumption (Joules)
(b) Deadline Violation Time (ms)
(c) Response Time (ms)
@@ III Results
This section presents an appraisal of the RIGEO approach’s performance through various metrics. The experimental configuration utilizes a laptop with an Intel Core i7 processor at 2.80 GHz, 16 GB RAM, and Windows 10 64-bit. Algorithms were implemented as simulations in MATLAB 2016, leveraging its computational capabilities to model the fog computing scenario, running 50 times using parallel processors.
@@
The study simulated a fog environment with interconnected nodes having diverse energy capacities and consumption rates. Tasks from IoT devices (200-600) were distributed among nodes with 2000-6000 MIPS processing power and 80-200 Joules energy usage. The traffic threshold was set at the median of observed traffic, and the deadline threshold at the 25th percentile of task deadlines, both recalculated periodically.
@@ III-B Deadline Violation
At 600 tasks, RIGEO demonstrated approximately 19% improvement over WCLA+GA and 14% improvement over ETFC in terms of deadline violation time. This capability proves especially beneficial in real-world scenarios where adherence to deadlines is critical for optimal performance and user satisfaction. RIGEO shows promise in enhancing system efficiency by minimizing deadline violations, evaluated using (3).

### v5

@@ I Introduction
This work proposes the Reinforcement Improved Golden Eagle Optimization (RIGEO) algorithm, an adaptive task scheduling framework. RIGEO dynamically classifies Fog Nodes (FNs) into low-traffic and high-traffic categories based on real-time network conditions. Low-traffic nodes are characterized by average network traffic that falls below a predefined threshold, whereas high-traffic nodes exceed this average. It employs Improved Golden Eagle Optimization (IGEO) for short-deadline tasks on low-traffic nodes and Reinforcement Learning for long-deadline tasks on high-traffic nodes, leveraging the strengths of each method for different operational scenarios [10].
RIGEO operates within a three-layered architecture: cloud, fog, and IoT devices. Figure 1 illustrates this hierarchy, where the fog layer comprises a controller and a network for resource management and task scheduling, while the cloud offers significant computing and storage capabilities.
@@ I-A Energy Efficiency
Vispute et al. [11] propose a PSO-based task scheduling scheme that outperforms existing approaches in makespan, energy consumption, and execution time.
@@ I-B Deadline Constraints
Azami et al. [1] incorporate two semi-greedy algorithms to minimize deadline violations for IoT tasks while ensuring QoS compliance, achieving a significant reduction in deadline violation time.
@@ I-C Latency Reduction
Klatoun et al. [7] propose a task scheduling framework for IoMT that processes data closer to end-users, reducing latency and enhancing response times.
Hosseini et al. [6] present an algorithm using fuzzy logic and AHP to prioritize tasks based on completion time, energy, RAM, and deadlines, effectively balancing cost and latency in mobile fog environments.
@@
Energy Deadline Latency
@@ II-A Problem Statement
Each task $t_{i}$ is described by a tuple $(c_{i},m_{i},d_{i},a_{i},q_{i})$ sampled as $c_{i}\sim\mathcal{U}\{1,\ldots,5\}$, $m_{i},d_{i},q_{i}\sim\mathcal{U}[0,1]$, and $a_{i}\sim\mathcal{U}\{1,\ldots,T\}$ where $T=10$, so batch sizes vary naturally across time slots. Each task also carries an implicit data payload whose size is proportional to $c_{i}$ (typically $s_{i}\approx 0.1$$2.5$ MB for IoT sensor and actuation data).
Each FN $f_{j}$ is described by a tuple $(C_{j},M_{j},\kappa_{j},\theta_{j},\varepsilon_{j},h_{j})$ where $C_{j}\sim\mathcal{U}\{81,\ldots,100\}$, $M_{j}\sim\mathcal{U}\{5,\ldots,16\}$ GB, $\kappa_{j}\sim\mathcal{U}[0,1]$, $\theta_{j}\sim\mathcal{U}[0.5,1.5]$, $\varepsilon_{j}=u_{j}+v_{j}$ with $u_{j}\sim\mathcal{U}\{3,4,5\}$ and $v_{j}\sim\mathcal{U}[0,1]$, and fixed $h_{j}=0.003$.
The fog network is modeled as a fully connected graph $G=(V,E)$ with $|V|=m$ nodes on high-bandwidth wired links ($B_{jk}\geq 1$ Gbps, $\delta_{jk}<0.1$ ms), making communication overhead negligible relative to execution time (see Section II-A1). A symmetric communication-preference matrix $P\in[0,1]^{m\times m}$ with $P(i,i)=0$ and $P(i,j)=P(j,i)\sim\mathcal{U}[0,1]$ is refined online by the RL stage.
@@ II-A1 Response time
The response time $R_{i}$ for task $t_{i}$ assigned to FN $f_{j}$ is [9]:
$R_{i}=P_{i}+T_{i}+E_{i}$ (1)
where the three components are defined as follows.
Propagation delay – the signal travel time over the physical link between the source and the assigned FN:
$P_{i}=\delta_{\text{src},\,j}$ (2)
Transmission delay – the time to transfer the task’s data payload $s_{i}$ over the link bandwidth:
$T_{i}=\frac{s_{i}}{B_{\text{src},\,j}}$ (3)
Execution time – the CPU processing time on the assigned node:
$E_{i}=\frac{c_{i}}{C_{j}+1}$ (4)
Since $P_{i}+T_{i}\ll E_{i}$, the per-task response time simplifies to $R_{i}\approx E_{i}$, and the system-wide metric is the makespan $\max_{j}W_{j}$, where $W_{j}=\sum_{x_{k}=j}E_{k}$. Deadline violations are thus:
$DV_{i}=\max\!\bigl(0,\;E_{i}-d_{i}\bigr)$ (5)
@@ II-A2 Deadline violation
A deadline violation occurs when a task’s completion time exceeds its specified deadline [4].
@@ II-B Simulation Environment
We evaluate $n\in\{200,300,400,500,600\}$ tasks on $m=20$ heterogeneous FNs over $T=10$ time slots (Table II), averaged over 50 Monte Carlo runs. FNs whose mean traffic exceeds the network-wide median are labeled high-traffic ($F_{\text{high}}$); the deadline threshold is the 25th percentile of each workload. Population-based optimizers use $N_{pop}=50$ and $N_{iter}=2$ per batch.
@@
Network topology Full mesh
@@
FN CPU capacity ($C_{j}$) $\mathcal{U}\{81,\ldots,100\}$ units
FN memory ($M_{j}$) $\mathcal{U}\{5,\ldots,16\}$ GB
FN energy rate ($\varepsilon_{j}$) $\mathcal{U}\{3,4,5\}+\mathcal{U}[0,1)$ J/cycle
@@
FN cost rate ($\kappa_{j}$) $\mathcal{U}[0,1)$
@@
Task data size ($s_{i}$) $\leq c_{i}\times 0.5\approx 2.5$ MB
Link bandwidth ($B_{jk}$) $\geq 1$ Gbps (wired LAN)
Propagation delay ($\delta_{jk}$) <0.1<0.1 ms (co-located)
Communication overhead ($P_{i}+T_{i}$) Negligible ($\ll E_{i}$); see Sec. II-A1
Deadline threshold 25th percentile
Traffic threshold Median of node mean
@@
II-C Proposed Method
At the start of each experiment, traffic labels and the deadline threshold are computed and used to partition FNs and tasks.
When network traffic is high, faster algorithms must be used to ensure that tasks are executed within their deadlines.
@@ II-C1 Reinforcement Learning for Long-Deadline Tasks
Reward and penalty. After task $t_{i}$ is assigned to FN $a$, the response time $R_{i}\approx E_{i}=c_{i}/(C_{a}+1)$ (see Section II-A1) is compared with deadline $d_{i}$. A met deadline rewards the transition (scales up by 10%); a violation penalises it (scales down by 10%):
@@ II-C2 Golden Eagle Optimization (GEO)
GEO is a swarm-based meta-heuristic whose search mechanism balances exploration (searching new areas) and exploitation (refining solutions). It adjusts speed during different hunting stages to find global optima while avoiding local optima.
In GEO, each eagle selects a target, calculates its cruising vector, and circles the best location found. The prey represents the optimal solution, with each eagle tracking its own best discovery. The attack vector for eagle $i$ is calculated through Eq. (12), where $\vec{X}_{f}^{*}$ is the position vector of the best prey found so far and $\vec{X}_{i}$ is the current position vector of eagle $i$.
@@ II-C3 Improved Golden Eagle Optimization (IGEO)
Figure 2: Mutation operator. Green segments indicate genes subject to mutation, Blue segments remain constant.
@@
Figure 3: Crossover operations: (a) one-point and (b) two-point. Purple and Green indicate exchanged segments.
@@
0:  Task $T$ with deadline $D$, Fog nodes $F=\{f_{1},\ldots,f_{n}\}$
@@
1:  $Low\_Traffic\_FNs\leftarrow\emptyset$, $High\_Traffic\_FNs\leftarrow\emptyset$
2:  for each $f\in F$ do
3:   if f.traffic<traffic_thresholdf.traffic<traffic\_threshold then
4:    Add $f$ to Low_Traffic_FNsLow\_Traffic\_FNs
5:   else
6:    Add $f$ to High_Traffic_FNsHigh\_Traffic\_FNs
7:   end if
8:  end for
9:  if T.deadline<deadline_thresholdT.deadline<deadline\_threshold then
10:   $f\leftarrow SELECT(Low\_Traffic\_FNs)$
11:   $result\leftarrow IGEO\_PROCESS(T,f)$
12:  else
13:   $f\leftarrow SELECT(High\_Traffic\_FNs)$
14:   $result\leftarrow RL\_PROCESS(T,f)$
15:  end if
16:  return resultresult
@@ III Results
This section presents an appraisal of the RIGEO approach’s performance through various metrics. Algorithms were implemented in MATLAB 2016 on an Intel Core i7 2.80 GHz machine with 16 GB RAM.
@@
Simulation parameters are detailed in Table II.
@@ III-B Deadline Violation
At 600 tasks, RIGEO demonstrated approximately 19% improvement over WCLA+GA and 14% improvement over ETFC in terms of deadline violation time.